OpenStudy (anonymous):

if 5 is a root of the equation x^2- kx-10 =0 , what is the value of k?

jimthompson5910 (jim_thompson5910):

If 5 is a root, then this means that x = 5 would make the entire left side equal to 0. So simply plug in x = 5 and solve for k. x^2- kx-10 =0 (5)^2- k(5)-10 =0 25 - 5k - 10 = 0 15 - 5k = 0 -5k = -15 k = -15/(-5) k = 3

OpenStudy (cwrw238):

( x-5)(x+2) = x^2 - 3x - 10 = 0 k = 3